OpenStudy (oswaldmurphy):

A particle is moving along a projectile path where the initial height is 160 feet with an initial speed of 144 feet per second. What is the maximum height of the particle? 128 feet 224 feet 272 feet 484 feet

OpenStudy (oswaldmurphy):

First thing I did was get the formula for the gravity. s(t) = -16t^2 + 144t + 160 I plugged it in. What do I do next?

OpenStudy (oswaldmurphy):

@Jadeishere

OpenStudy (jadeishere):

The formula would be s(t) = -16t^2 +vt + s Plug in the values; s(t) = -16t^2 + 144t + 160

OpenStudy (oswaldmurphy):

What next?

OpenStudy (jadeishere):

I'm not quite sure.. Like I said, not the best with math. Anyone would be better than me. @agent0smith @AloneS @imqwerty @SIenderman

OpenStudy (jadeishere):

@welshfella

OpenStudy (oswaldmurphy):

@sammixboo

OpenStudy (welshfella):

find the derivative ds /dt This will give you an expression for the velocity. When the projectile is at its maximum height the value of v will be 0. v = ds/dt = -32t + 144 = 0 t = 144/32 = 4.5 seconds

OpenStudy (welshfella):

Now to find maximum height s(4.5) = -16(4.5)^2 + 144(4.5) + 160

OpenStudy (welshfella):

grab ur calculator and work this out

OpenStudy (welshfella):

have you done any calculus? derivatives?

OpenStudy (oswaldmurphy):

No

OpenStudy (welshfella):

If not then we can use another constant acceleration question ( which came form calculus anyway) v = u + at u = initial speed = 144, a = acceleration due to gravity = -32 , and v = 0 9 at maximum height) 0 = 144 - 32t - giving t = 4.5 secs

OpenStudy (oswaldmurphy):

I just dont understand how to get t?

OpenStudy (oswaldmurphy):

@welshfella

OpenStudy (welshfella):

0 = 144 - 32t solve for t subtract 144 from both sides -144 = -32t now divide both sides by -32

OpenStudy (oswaldmurphy):

Where is 32 from? Is that a second problem?

OpenStudy (oswaldmurphy):

to solve for t its 0 = v - st ?

OpenStudy (faiqraees):

You need help?

OpenStudy (oswaldmurphy):

Yes @FaiqRaees

OpenStudy (faiqraees):

First I am going to clear the assumptions we're taking in this question(Since the question is vague in every aspect) 1. The object has initial velocity in the vertical direction 2. Gravitational acceleration is opposite to the initial velocity 3. da/dh = 0 and acceleration =9.81 Clear?

OpenStudy (oswaldmurphy):

My tacher said do -b/2a.... like aos

OpenStudy (faiqraees):

We're going to suppose the initial height is 0. (Meaning our ground has risen to 160 feet) Use the formula \[\large\rm (V_{final})^2=(V_{initial})^2-2(acceleration)(distance)\ \] \[ \color{red}{\text{ Note:- The sign after initial is changed because acceleration is negative.}} \]

OpenStudy (oswaldmurphy):

When I do that I get 4.5. s(4.5) = -16(4.5)^2 + 144(4.5) + 160 When I do that I get 252...

OpenStudy (faiqraees):

What values are you taking for each variable?

OpenStudy (oswaldmurphy):

t = 4.5 a: -16 b: 144

OpenStudy (oswaldmurphy):

-144/-32 is 4.5.

OpenStudy (faiqraees):

Which formula are you using?

OpenStudy (oswaldmurphy):

-b/2a

OpenStudy (faiqraees):

That's a wrong formula, use the one I provided

OpenStudy (oswaldmurphy):

WHen I am done I get 484. Is that the answer?

OpenStudy (faiqraees):

Correct

OpenStudy (oswaldmurphy):

What I did was WAY easier.. XD I did simply -b/2a and got t. Then I solved for s(t) and got it :P thank you!

OpenStudy (faiqraees):

Good

OpenStudy (faiqraees):

But in my method you just had to plug the respective values and then add 160

OpenStudy (oswaldmurphy):

I havent learned that yet.. I dont even know what a respective value it. This is Algebra 1 and im 13 XD

OpenStudy (faiqraees):

Well then you should next time specify that. In this way the others will keep the proficiency level of their explanation to minimal

OpenStudy (oswaldmurphy):

Sorry D:

OpenStudy (agent0smith):

lol poor guy first got calculus help, then physics help. Probably a good idea to next time specify that you're in algebra 1 @OswaldMurphy since these type of questions don't make it clear what type of class you're in.

OpenStudy (jadeishere):

Super sorry :( I was busy doing other stuff and I should've stayed here to help you... @OswaldMurphy

OpenStudy (oswaldmurphy):

@agent0smith I didnt know this question had different solutions. Thanks @Jadeishere

OpenStudy (oswaldmurphy):

Its ok @Jadeishere *

OpenStudy (agent0smith):

Well it can be done as a physics problem, an algebra problem, or a calculus problem. They'll all get the same answer, but they're all pretty different methods.